# Multi-scale Autoregressive Models are Laplacian, Discrete, and Latent Diffusion Models in Disguise

### v1

@@ 6 Limitations and Future Work
Our empirical study is intentionally small: we use controlled MLP baselines to isolate effects under limited compute. A more complete assessment would train a reference VAR model end to end on standard image benchmarks and compare to its diffusion analogue under similar settings. Specifically: (i) latent vs. pixel: run next-scale refinement in VQ-VAE latents and, as a “no-latent” control, in a Laplacian pixel pyramid; (ii) discrete vs. continuous: keep the pyramid but replace the categorical head with a Gaussian regression head to test quantisation; (iii) multi-scale vs. single-scale: collapse to a single latent grid to assess the value of coarse to fine. For each setting, attach a diffusion head that predicts $\varepsilon$ with $T$ steps in the same space (pixel diffusion, latent diffusion, and discrete diffusion, respectively). Report quality–cost frontiers (e.g., FID or PSNR versus wall-clock and versus network evaluations), calibration (e.g., bits/dim with standard dequantization), and sensitivity to $S$ scales versus $T$ steps.
@@
In vanilla VAR, the forward path that builds the latent/residual pyramid is fully deterministic. Stochasticity appears only in bitwise self-correction variants (e.g., [15]) that inject Bernoulli noise at the code level. Whether mild forward randomness helps remains an open question and requires further experiments.
Leverage the diffusion ecosystem.
Several techniques from diffusion may transfer directly to VAR’s iterative-refinement view. Examples include classifier-free guidance adapted to categorical heads, self-conditioning across scales, loss re-weighting akin to simplified diffusion objectives, few-step distillation of a large-$S$ refiner into a small-$S$ student, early-exit or schedule search analogous to noise-schedule tuning, and solver-style improvements such as consistency training to reduce the number of refinement transitions. One might also explore rectified-flow style objectives between consecutive scales, latent-consistency adapters for faster sampling, and temperature or top-$k$ calibration per scale to trade fidelity for diversity. Given the fruitful research in the diffusion literature, it is now the time for VAR to leverage many techniques from this field.
@@ 7 Conclusion
We reframed Visual Autoregressive models as iterative refiners operating on a latent Laplacian pyramid. This view makes the generation path explicit: a short sequence of coarse-to-fine residual updates, and clarifies how VAR relates to denoising diffusion while retaining exact cross-scale factorisation and within-scale parallelism. Under this lens, three design choices appear central: working in a compact latent space, predicting discrete indices rather than continuous targets, and allocating supervision across frequency bands. Small toy on MNIST support these hypotheses, indicating that a modest number of banded steps can capture most of the benefit.

### v2

@@ 6 Limitations and Future Work
Our empirical study is intentionally limited in scope. We use controlled MLP baselines to isolate specific factors under modest compute budgets. A broader assessment should train a full VAR model end-to-end on standard image benchmarks and compare it with a diffusion baseline under matched architectures and training budgets. Useful ablations include: (i) latent vs. pixel, by running next-scale refinement in VQ-VAE latent space and in a Laplacian pixel pyramid as a no-latent control; (ii) discrete vs. continuous, by replacing categorical prediction heads with Gaussian regression heads while keeping the same pyramid structure; and (iii) multi-scale vs. single-scale, by collapsing to a single latent grid. For each setting, one can add a diffusion objective in the same representation space (pixel, latent, or discrete) and compare quality–cost trade-offs (e.g., FID/PSNR versus wall-clock and function evaluations), calibration (e.g., bits/dim with standard dequantisation), and sensitivity to refinement depth $S$ versus diffusion steps $T$.
@@
In standard VAR, the forward construction of the latent/residual pyramid is deterministic. Stochasticity is introduced only in some variants, such as bitwise self-correction methods that inject Bernoulli noise at the code level [10]. The effect of adding controlled stochasticity to the forward process is not yet clear and warrants systematic evaluation.
Leveraging diffusion methods.
The iterative-refinement view suggests several diffusion-inspired techniques that may be applicable to VAR. Candidate directions include classifier-free guidance for categorical heads, self-conditioning across scales, loss reweighting, few-step distillation from larger-$S$ models, and schedule design or early-exit strategies. Other possibilities include consistency-style objectives between scales, latent-consistency adapters for faster sampling, and per-scale temperature or top-$k$ calibration to tune fidelity-diversity trade-offs. These directions should be treated as hypotheses for future empirical study.
@@ 7 Conclusion
We reframed Visual Autoregressive models as iterative refiners operating on a latent Laplacian pyramid. This view makes the generation path explicit as a short sequence of coarse-to-fine residual updates and clarifies how VAR relates to denoising diffusion while retaining exact cross-scale factorisation and within-scale parallelism. Under this lens, three design choices appear central: working in a compact latent space, predicting discrete indices rather than continuous targets, and allocating supervision across frequency bands. Small MNIST toy experiments support these hypotheses, indicating that a modest number of banded steps can capture most of the benefit.